# Physical meaning of non-extensive term in Massieu functions

### v1

@@ III.2 General derivations
$\tilde{\Psi}(\hat{\vec{X}}^{s},\vec{Y}^{r-s})=\ln P(\hat{\vec{X}}^{s})+\tilde{\Psi}(\vec{Y})+Y_{1}\hat{X}_{1}+\dots+Y_{s}\hat{X}_{s}.$ (30)
@@
$\left\langle\frac{\partial\tilde{\Psi}(\hat{\vec{X}}^{s},\vec{Y}^{r-s})}{\partial\hat{X}_{i}}\right\rangle$ $=Y_{i}+\left\langle\frac{1}{P(\hat{\vec{X}}^{s})}\frac{\partial P(\hat{\vec{X}}^{s})}{\partial\hat{X}_{i}}\right\rangle$
@@
$\left\langle\frac{\partial\tilde{\Psi}(\hat{\vec{X}}^{s},\vec{Y}^{r-s})}{\partial\hat{X}_{i}}\right\rangle=\frac{\partial\Psi(\vec{X}^{s},\vec{Y}^{r-s})}{\partial X_{i}}.$ (32)
@@ A Taylor expansion around $\vec{X}^{s}$ is used to evaluate the average:
$\left\langle\frac{\partial\tilde{\Psi}(\hat{\vec{X}}^{s},\vec{Y}^{r-s})}{\partial\hat{X}_{i}}\right\rangle$ $=\frac{\partial\tilde{\Psi}(\vec{X}^{s},\vec{Y}^{r-s})}{\partial X_{i}}+$
$\frac{1}{2}\sum_{j,k}\frac{\partial^{3}\Psi(\vec{X}^{s},\vec{Y}^{r-s})}{\partial X_{i}\partial X_{j}\partial X_{k}}\langle\Delta\hat{X}_{j}\Delta\hat{X}_{k}\rangle,$ (33)
with $\Delta\hat{X}_{j}=\hat{X}_{j}-X_{j}$ and terms $O(N^{-2})$ are neglected (including the difference between $\Psi$ and $\tilde{\Psi}$ in the third derivative). Therefore,
$\frac{\partial}{\partial X_{i}}$ $[\tilde{\Psi}(\vec{X}^{s},\vec{Y}^{r-s})-\Psi(\vec{X}^{s},\vec{Y}^{r-s})]$
$=-\frac{1}{2}\sum_{j,k}\frac{\partial^{3}\Psi(\vec{X}^{s},\vec{Y}^{r-s})}{\partial X_{i}\partial X_{j}\partial X_{k}}\langle\Delta\hat{X}_{j}\Delta\hat{X}_{k}\rangle$
@@
$\frac{\partial\Delta\tilde{\Psi}(\vec{X}^{s},\vec{Y}^{r-s})}{\partial X_{i}}$ $=\frac{1}{2}\sum_{j,k}\frac{\partial(C^{-1})_{jk}}{\partial X_{i}}C_{jk}$
@@
where we define $\Delta\tilde{\Psi}(\vec{X}^{s},\vec{Y}^{r-s})\equiv\tilde{\Psi}(\vec{X}^{s},\vec{Y}^{r-s})-\Psi(\vec{X}^{s},\vec{Y}^{r-s})$. It can be shown (see Sec. 15.8 in Harville (1997)) that the solution of this equation system is
$\Delta\tilde{\Psi}(\vec{X}^{s},\vec{Y}^{r-s})=-\frac{1}{2}\ln(\det C)+c,$ (37)
@@
$\Delta\tilde{\Psi}_{\text{id}}(\vec{X}^{s},\vec{Y}^{r-s})=-\frac{1}{2}\ln(\det C_{\text{id}})+c.$ (38)
@@
$\Delta\tilde{\Psi}_{\text{ex}}(\vec{X}^{s},\vec{Y}^{r-s})$ $=\tilde{\Psi}_{\text{ex}}(\vec{X}^{s},\vec{Y}^{r-s})-\Psi_{\text{ex}}(\vec{X}^{s},\vec{Y}^{r-s})$
@@
Using the result of Eq. (37), and the normalization of the probability distribution of fluctuations, it can be shown that the difference $\Delta\tilde{\Psi}(\vec{Y})$ for the last Legendre transform is a constant; see Appendix A. It is known that this constant is zero for some solvable systems, meaning that $\tilde{\Psi}(\vec{Y})$ is exactly extensive; this is the case for the one-dimensional spin lattice, as is mentioned in Sec. V.2 below, and for the grand canonical ensemble of an ideal gas (see Sec. 8.2 in Swendsen (2017), where the Massieu function that depends on $\vec{Y}=(\frac{1}{T},\frac{\mu}{T})$ is called the grand canonical entropy).
@@ IV.1 Ideal case
$\tilde{\Psi}_{\text{id}}(\tfrac{1}{T},N)=\ln\mathcal{Z}_{\text{id}}=N\ln\left(\frac{V}{\lambda^{3}N}\right)+N-\frac{1}{2}\ln N$ (40)
@@
$\Psi_{\text{id}}(\tfrac{1}{T},N)=N\ln\left(\frac{V}{\lambda^{3}N}\right)+N$ (41)
The difference $\tilde{\Psi}_{\text{id}}(\tfrac{1}{T},N)-\Psi_{\text{id}}(\tfrac{1}{T},N)$ is
$\Delta\tilde{\Psi}_{\text{id}}(\tfrac{1}{T},N)=-\frac{1}{2}\ln N.$ (42)
@@ IV.2 Non ideal case
As before, we analyze the system of particles described by $\hat{U}$, $\hat{N}$ and $V$, where energy and particle number fluctuate and volume is constant. We are interested in the non-extensive term of the Massieu function that corresponds to the excess free energy: $\tilde{\Psi}_{\text{ex}}(\tfrac{1}{T},N)=-\tilde{F}_{\text{ex}}/T$. This quantity is related to the particle fluctuations as can be seen in Eq. (29). The mean squared fluctuations are [see Eq. (24)]
@@
$\Delta\tilde{\Psi}_{\text{ex}}(\tfrac{1}{T},N)=\frac{1}{2}\ln\Gamma,$ (44)
@@
Next, we contrast the analytical expression against numerical results for hard disks and spheres. Let us consider first hard disk interaction. Solana et al. have proposed a simple equation of state (EOS) for a system of hard disks Mulero et al. (2009),
@@ V Spin lattice
Let us consider a $d$-dimensional lattice with $N$ spins; each spin has two possible states: up or down. There are $\hat{N}_{+}$ and $\hat{N}_{-}=N-\hat{N}_{+}$ spines in the up and down states. The total number of spins, $N$, is constant and plays the same role as the volume in the previous example. Each spin, $s_{i}$, takes the value 1 or $-1$. The microscopic state is given by $\vec{s}=(s_{1},\dots,s_{N})$ and its energy is
@@
$e^{\tilde{\Psi}(\hat{M},\hat{E})}$ $=\Omega(\hat{M},\hat{E})$ (57)
$e^{\tilde{\Psi}(\hat{M},\frac{1}{T})}$ $=\sum_{\hat{E}}e^{\tilde{\Psi}(\hat{M},\hat{E})}e^{-\frac{1}{T}\hat{E}}$ (58)
$e^{\tilde{\Psi}(\frac{h}{T},\frac{1}{T})}$ $=\sum_{\hat{M}}e^{\tilde{\Psi}(\hat{M},\frac{1}{T})}e^{\frac{h}{T}\hat{M}}.$ (59)
@@ V.1 Ideal case
$e^{\tilde{\Psi}_{\text{id}}(\frac{h}{T})}=\sum_{\hat{M}}e^{\tilde{\Psi}_{\text{id}}(\hat{M})}e^{\frac{h}{T}\hat{M}},$ (60)
@@
$\tilde{\Psi}_{\text{id}}(\hat{M})=\ln\Omega(\hat{M})=\ln\frac{N!}{\hat{N}_{+}!\hat{N}_{-}!},$ (61)
with $\Omega(\hat{M})$ the number of microstates with magnetization $\hat{M}$. Let us consider $\tilde{\Psi}_{\text{id}}(M)$, evaluated at the thermodynamic average $M$, and apply Stirling’s approximation (up to order $\ln N$):
$\tilde{\Psi}_{\text{id}}(M)=\underbrace{N_{+}\ln\frac{N}{N_{+}}+N_{-}\ln\frac{N}{N_{-}}}_{O(N)}-\underbrace{\frac{1}{2}\ln\frac{N_{+}N_{-}}{N}}_{O(\ln N)}.$ (62)
@@
$\Psi_{\text{id}}(M)=N_{+}\ln\frac{N}{N_{+}}+N_{-}\ln\frac{N}{N_{-}}.$ (63)
@@
$\Delta\tilde{\Psi}_{\text{id}}(M)=-\frac{1}{2}\ln\frac{N_{+}N_{-}}{N}.$ (64)
@@
$\Delta\tilde{\Psi}_{\text{id}}(M)=-\frac{1}{2}\ln\langle(\Delta\hat{M})^{2}\rangle_{\text{id}},$ (65)
@@
$\langle(\Delta\hat{M})^{2}\rangle_{\text{id}}=\frac{\partial M}{\partial\frac{h}{T}}=\frac{-1}{\frac{\partial^{2}\Psi_{\text{id}}(M)}{\partial M^{2}}}=4\frac{N_{+}N_{-}}{N},$ (66)
where we have used that $\frac{h}{T}=-\frac{\partial\Psi_{\text{id}}(M)}{\partial M}$ (relationship that can be obtained from the Legendre transform $\Psi_{\text{id}}(\frac{h}{T})=\Psi_{\text{id}}(M)+\frac{h}{T}M$). Therefore, comparing Eqs. (64) and (65) it can be seen that the theory correctly predicts the non-extensive term of the Massieu function at order $\ln N$.
@@ V.2 Non ideal case
$\Psi(M,\tfrac{1}{T})=\Psi(\tfrac{h}{T},\tfrac{1}{T})-\tfrac{h}{T}M.$ (67)
$\Psi(M,\tfrac{1}{T})$ and $\Psi(\tfrac{h}{T},\tfrac{1}{T})$ are associated to the Helmholtz and Gibbs free energies of the spin lattice, respectively. Let us consider the exact solution of the Ising model of $N$ spins in 1D with periodic boundary conditions; $\ln\mathcal{Z}(\tfrac{h}{T},\tfrac{1}{T})$ is extensive, so $\Psi(\tfrac{h}{T},\tfrac{1}{T})=\tilde{\Psi}(\tfrac{h}{T},\tfrac{1}{T})$; see, for example, Sec. 5.6.1 in Reichl (1998) or Sec. 6.5.2 in Schwabl (2006). The result is
$\Psi($ $\tfrac{h}{T},\tfrac{1}{T})=N\frac{\epsilon}{T}$
@@
From $M=\frac{\partial\Psi(h/T,1/T)}{\partial(h/T)}$ we can obtain $\tfrac{h}{T}$ as a function of $M$ and $\tfrac{1}{T}$, and using this solution in Eq. (67) we have $\Psi(M,\frac{1}{T})$. The detailed expressions are given in Appendix C. We wish to verify Eq. (39) that, in the present situation, reads
$\Delta\tilde{\Psi}_{\text{ex}}(M,\tfrac{1}{T})$ $=\tilde{\Psi}_{\text{ex}}(M,\tfrac{1}{T})-\Psi_{\text{ex}}(M,\tfrac{1}{T})$
@@
where $\Psi_{\text{ex}}(M,\frac{1}{T})=\Psi(M,\frac{1}{T})-\Psi_{\text{id}}(M)$, and $\Psi_{\text{id}}(M)$ is given by (63). The mean squared fluctuations are $\langle(\Delta\hat{M})^{2}\rangle=-(\frac{\partial^{2}\Psi(M,1/T)}{\partial M^{2}})^{-1}$; the ideal fluctuations were already calculated in (66).
We calculate $\tilde{\Psi}_{\text{ex}}(M,\frac{1}{T})=\tilde{\Psi}(M,\frac{1}{T})-\tilde{\Psi}_{\text{id}}(M)$ numerically. From (58), evaluated at $\hat{M}=M$, we have
$\tilde{\Psi}(M,\tfrac{1}{T})=\ln\left(\sum_{\hat{E}}\Omega(M,\hat{E})e^{-\frac{1}{T}\hat{E}}\right)=\ln\sum_{\omega_{M}}e^{-\frac{1}{T}\hat{E}}.$ (70)
In order to evaluate the sum in the last equation we consider a system small enough so that all possible microstates $\omega_{M}$, for a given magnetization $M$, can be computed. The ideal value, $\tilde{\Psi}_{\text{id}}(M)$, is given by Eq. (61) with $\hat{M}=M$. The numerical results are obtained for values of $\epsilon/T$ between 0 and 1.2, and for $M=N/2$. Figure 3 shows results for $N$ equal 16 and 32. As the system size increases, numerical results approach the theoretical prediction of Eq. (V.2). This is an expected behavior since the theoretical results hold for large $N$.
Figure 3: Non extensive term $\tilde{\Psi}_{\text{ex}}(M,\tfrac{1}{T})-\Psi_{\text{ex}}(M,\tfrac{1}{T})$ against $\epsilon/T$ for the one-dimensional spin lattice. The curve corresponds to Eq. (V.2). Dots where numerically obtained for $N=16$ and 32; in both cases, $M=N/2$. Numerical results approach the theoretical curve as $N$ increases.

### v2

@@ III.2 General derivations
$\tilde{S}[\hat{\vec{X}}^{s},\vec{Y}^{r-s}]=\ln P(\hat{\vec{X}}^{s})+\tilde{S}[\vec{Y}]+Y_{1}\hat{X}_{1}+\dots+Y_{s}\hat{X}_{s}.$ (30)
@@
$\left\langle\frac{\partial\tilde{S}[\hat{\vec{X}}^{s},\vec{Y}^{r-s}]}{\partial\hat{X}_{i}}\right\rangle$ $=Y_{i}+\left\langle\frac{1}{P(\hat{\vec{X}}^{s})}\frac{\partial P(\hat{\vec{X}}^{s})}{\partial\hat{X}_{i}}\right\rangle$
@@
$\left\langle\frac{\partial\tilde{S}[\hat{\vec{X}}^{s},\vec{Y}^{r-s}]}{\partial\hat{X}_{i}}\right\rangle=\frac{\partial S[\vec{X}^{s},\vec{Y}^{r-s}]}{\partial X_{i}}.$ (32)
@@ A Taylor expansion around $\vec{X}^{s}$ is used to evaluate the average:
$\left\langle\frac{\partial\tilde{S}[\hat{\vec{X}}^{s},\vec{Y}^{r-s}]}{\partial\hat{X}_{i}}\right\rangle$ $=\frac{\partial\tilde{S}[\vec{X}^{s},\vec{Y}^{r-s}]}{\partial X_{i}}+$
$\frac{1}{2}\sum_{j,k}\frac{\partial^{3}S[\vec{X}^{s},\vec{Y}^{r-s}]}{\partial X_{i}\partial X_{j}\partial X_{k}}\langle\Delta\hat{X}_{j}\Delta\hat{X}_{k}\rangle,$ (33)
with $\Delta\hat{X}_{j}=\hat{X}_{j}-X_{j}$ and terms $O(N^{-2})$ are neglected (including the difference between $S$ and $\tilde{S}$ in the third derivative). Therefore,
$\frac{\partial}{\partial X_{i}}$ $(\tilde{S}[\vec{X}^{s},\vec{Y}^{r-s}]-S[\vec{X}^{s},\vec{Y}^{r-s}])$
$=-\frac{1}{2}\sum_{j,k}\frac{\partial^{3}S[\vec{X}^{s},\vec{Y}^{r-s}]}{\partial X_{i}\partial X_{j}\partial X_{k}}\langle\Delta\hat{X}_{j}\Delta\hat{X}_{k}\rangle$
@@
$\frac{\partial\Delta\tilde{S}[\vec{X}^{s},\vec{Y}^{r-s}]}{\partial X_{i}}$ $=\frac{1}{2}\sum_{j,k}\frac{\partial(C^{-1})_{jk}}{\partial X_{i}}C_{jk}$
@@
where we define $\Delta\tilde{S}[\vec{X}^{s},\vec{Y}^{r-s}]\equiv\tilde{S}[\vec{X}^{s},\vec{Y}^{r-s}]-S[\vec{X}^{s},\vec{Y}^{r-s}]$. It can be shown (see Sec. 15.8 in Harville (1997)) that the solution of this equation system is
$\Delta\tilde{S}[\vec{X}^{s},\vec{Y}^{r-s}]=-\frac{1}{2}\ln(\det C)+c,$ (37)
@@
$\Delta\tilde{S}_{\text{id}}[\vec{X}^{s},\vec{Y}^{r-s}]=-\frac{1}{2}\ln(\det C_{\text{id}})+c.$ (38)
@@
$\Delta\tilde{S}_{\text{ex}}[\vec{X}^{s},\vec{Y}^{r-s}]$ $=\tilde{S}_{\text{ex}}[\vec{X}^{s},\vec{Y}^{r-s}]-S_{\text{ex}}[\vec{X}^{s},\vec{Y}^{r-s}]$
@@
Using the result of Eq. (37), and the normalization of the probability distribution of fluctuations, it can be shown that the difference $\Delta\tilde{S}[\vec{Y}]$ for the last Legendre transform is a constant; see Appendix A. It is known that this constant is zero for some solvable systems, meaning that $\tilde{S}[\vec{Y}]$ is exactly extensive; this is the case for the one-dimensional spin lattice, as is mentioned in Sec. V.2 below, and for the grand canonical ensemble of an ideal gas (see Sec. 8.2 in Swendsen (2017), where the Massieu function that depends on $\vec{Y}=(\frac{1}{T},\frac{\mu}{T})$ is called the grand canonical entropy).
@@ IV.1 Ideal case
$\tilde{S}_{\text{id}}[\tfrac{1}{T},N]=\ln\mathcal{Z}_{\text{id}}=N\ln\left(\frac{V}{\lambda^{3}N}\right)+N-\frac{1}{2}\ln N$ (40)
@@
$S_{\text{id}}[\tfrac{1}{T},N]=N\ln\left(\frac{V}{\lambda^{3}N}\right)+N$ (41)
The difference $\tilde{S}_{\text{id}}[\tfrac{1}{T},N]-S_{\text{id}}[\tfrac{1}{T},N]$ is
$\Delta\tilde{S}_{\text{id}}[\tfrac{1}{T},N]=-\frac{1}{2}\ln N.$ (42)
@@ IV.2 Non ideal case
As before, we analyze the system of particles described by $\hat{U}$, $\hat{N}$ and $V$, where energy and particle number fluctuate and volume is constant. We are interested in the non-extensive term of the Massieu function that corresponds to the excess free energy: $\tilde{S}_{\text{ex}}[\tfrac{1}{T},N]=-\tilde{F}_{\text{ex}}/T$. This quantity is related to the particle fluctuations as can be seen in Eq. (29). The mean squared fluctuations are [see Eq. (24)]
@@
$\Delta\tilde{S}_{\text{ex}}[\tfrac{1}{T},N]=\frac{1}{2}\ln\Gamma,$ (44)
@@
Next, we contrast the analytical expression against numerical results for hard disks and spheres. Let us consider first hard disk interaction. Mulero et al. have proposed a simple equation of state (EOS) for a system of hard disks Mulero et al. (2009),
@@ V Spin lattice
Let us consider a $d$-dimensional lattice with $N$ spins; each spin has two possible states: up or down. There are $\hat{N}_{+}$ and $\hat{N}_{-}=N-\hat{N}_{+}$ spins in the up and down states. The total number of spins, $N$, is constant and plays the same role as the volume in the previous example. Each spin, $s_{i}$, takes the value 1 or $-1$. The microscopic state is given by $\vec{s}=(s_{1},\dots,s_{N})$ and its energy is
@@
$e^{\tilde{S}(\hat{M},\hat{E})}$ $=\Omega(\hat{M},\hat{E})$ (57)
$e^{\tilde{S}[\hat{M},\frac{1}{T}]}$ $=\sum_{\hat{E}}e^{\tilde{S}(\hat{M},\hat{E})}e^{-\frac{1}{T}\hat{E}}$ (58)
$e^{\tilde{S}[\frac{h}{T},\frac{1}{T}]}$ $=\sum_{\hat{M}}e^{\tilde{S}[\hat{M},\frac{1}{T}]}e^{\frac{h}{T}\hat{M}}.$ (59)
@@ V.1 Ideal case
$e^{\tilde{S}_{\text{id}}[\frac{h}{T}]}=\sum_{\hat{M}}e^{\tilde{S}_{\text{id}}(\hat{M})}e^{\frac{h}{T}\hat{M}},$ (60)
@@
$\tilde{S}_{\text{id}}(\hat{M})=\ln\Omega(\hat{M})=\ln\frac{N!}{\hat{N}_{+}!\hat{N}_{-}!},$ (61)
with $\Omega(\hat{M})$ the number of microstates with magnetization $\hat{M}$. Let us consider $\tilde{S}_{\text{id}}(M)$, evaluated at the thermodynamic average $M$, and apply Stirling’s approximation (up to order $\ln N$):
$\tilde{S}_{\text{id}}(M)=\underbrace{N_{+}\ln\frac{N}{N_{+}}+N_{-}\ln\frac{N}{N_{-}}}_{O(N)}-\underbrace{\frac{1}{2}\ln\frac{N_{+}N_{-}}{N}}_{O(\ln N)}.$ (62)
@@
$S_{\text{id}}(M)=N_{+}\ln\frac{N}{N_{+}}+N_{-}\ln\frac{N}{N_{-}}.$ (63)
@@
$\Delta\tilde{S}_{\text{id}}(M)=-\frac{1}{2}\ln\frac{N_{+}N_{-}}{N}.$ (64)
@@
$\Delta\tilde{S}_{\text{id}}(M)=-\frac{1}{2}\ln\langle(\Delta\hat{M})^{2}\rangle_{\text{id}},$ (65)
@@
$\langle(\Delta\hat{M})^{2}\rangle_{\text{id}}=\frac{\partial M}{\partial\frac{h}{T}}=\frac{-1}{\frac{\partial^{2}S_{\text{id}}(M)}{\partial M^{2}}}=4\frac{N_{+}N_{-}}{N},$ (66)
where we have used that $\frac{h}{T}=-\frac{\partial S_{\text{id}}(M)}{\partial M}$ (relationship that can be obtained from the Legendre transform $S_{\text{id}}[\frac{h}{T}]=S_{\text{id}}(M)+\frac{h}{T}M$). Therefore, comparing Eqs. (64) and (65) it can be seen that the theory correctly predicts the non-extensive term of the Massieu function at order $\ln N$.
@@ V.2 Non ideal case
$S[M,\tfrac{1}{T}]=S[\tfrac{h}{T},\tfrac{1}{T}]-\tfrac{h}{T}M.$ (67)
$S[M,\tfrac{1}{T}]$ and $S[\tfrac{h}{T},\tfrac{1}{T}]$ are associated to the Helmholtz and Gibbs free energies of the spin lattice, respectively. Let us consider the exact solution of the Ising model of $N$ spins in 1D with periodic boundary conditions; $\ln\mathcal{Z}(\tfrac{h}{T},\tfrac{1}{T})$ is extensive, so $S[\tfrac{h}{T},\tfrac{1}{T}]=\tilde{S}[\tfrac{h}{T},\tfrac{1}{T}]$; see, for example, Sec. 5.6.1 in Reichl (1998) or Sec. 6.5.2 in Schwabl (2006). The result is
$S($ $\tfrac{h}{T},\tfrac{1}{T})=N\frac{\epsilon}{T}$
@@
From $M=\frac{\partial S[h/T,1/T]}{\partial(h/T)}$ we can obtain $\tfrac{h}{T}$ as a function of $M$ and $\tfrac{1}{T}$, and using this solution in Eq. (67) we have $S[M,\frac{1}{T}]$. The detailed expressions are given in Appendix C. We wish to verify Eq. (39) that, in the present situation, reads
$\Delta\tilde{S}_{\text{ex}}[M,\tfrac{1}{T}]$ $=\tilde{S}_{\text{ex}}[M,\tfrac{1}{T}]-S_{\text{ex}}[M,\tfrac{1}{T}]$
@@
where $S_{\text{ex}}[M,\frac{1}{T}]=S[M,\frac{1}{T}]-S_{\text{id}}(M)$, and $S_{\text{id}}(M)$ is given by (63). The mean squared fluctuations are $\langle(\Delta\hat{M})^{2}\rangle=-(\frac{\partial^{2}S[M,1/T]}{\partial M^{2}})^{-1}$; the ideal fluctuations were already calculated in (66).
We calculate $\tilde{S}_{\text{ex}}[M,\frac{1}{T}]=\tilde{S}[M,\frac{1}{T}]-\tilde{S}_{\text{id}}(M)$ numerically. From (58), evaluated at $\hat{M}=M$, we have
$\tilde{S}[M,\tfrac{1}{T}]=\ln\left(\sum_{\hat{E}}\Omega(M,\hat{E})e^{-\frac{1}{T}\hat{E}}\right)=\ln\sum_{\omega_{M}}e^{-\frac{1}{T}\hat{E}}.$ (70)
In order to evaluate the sum in the last equation we consider a system small enough so that all possible microstates $\omega_{M}$, for a given magnetization $M$, can be computed. The ideal value, $\tilde{S}_{\text{id}}(M)$, is given by Eq. (61) with $\hat{M}=M$. The numerical results are obtained for values of $\epsilon/T$ between 0 and 1.2, and for $M=N/2$. Figure 3 shows results for $N$ equal 16 and 32. As the system size increases, numerical results approach the theoretical prediction of Eq. (V.2). This is an expected behavior since the theoretical results hold for large $N$.
Figure 3: Non extensive term $\tilde{S}_{\text{ex}}[M,\tfrac{1}{T}]-S_{\text{ex}}[M,\tfrac{1}{T}]$ against $\epsilon/T$ for the one-dimensional spin lattice. The curve corresponds to Eq. (V.2). Dots where numerically obtained for $N=16$ and 32; in both cases, an external field is applied such that $M=N/2$. Numerical results approach the theoretical curve as $N$ increases.